# AdvReal: Adversarial Patch Generation Framework with Application to Adversarial Safety Evaluation of Object Detection Systems

### v1

@@ 4.3 Ablation Study
Tab. 4 presents ablation experiments that systematically examine the contributions of three modules in the AdvReal framework—namely, non-grid surfaces, Realistic Matching, and Shakedrop. The results reveal that each module significantly influences both the ASR and the perceptual quality AC of adversarial patches. Specifically, the exclusive use of the non-grid surfaces module boosts ASR from $87.45\%$ to $91.56\%$ while simultaneously lowering AC from $59.68\%$ to $55.64\%$, thereby enhancing both attack efficacy and visual quality. Similarly, when applied individually, both the Realistic Matching and Shakedrop modules improve ASR and reduce AC. Notably, the Realistic Matching module yields a particularly notable improvement, raising ASR to $90.48\%$ and reducing AC to $55.39\%$. Interestingly, the combined application of non-grid surfaces and Shakedrop results in a slight ASR decrease to $87.01\%$, yet it substantially elevates perceptual quality, as evidenced by an AC improvement to $54.28\%$, indicating a synergistic benefit. Ultimately, the full integration of all three modules delivers optimal performance, achieving an ASR of $93.72\%$ and an AC of $53.17\%$. These findings ultimately confirm the cumulative effectiveness of the proposed modules, underscoring their synergistic value in balancing attack potency with perceptual quality.
@@ 4.4 Robustness
When the adversarial patch is partially occluded, its attack efficacy is markedly reduced. To evaluate the robustness of different algorithms under occlusion, we superimpose a 100×100 pixel gray patch at the center of a 300×300 pixel adversarial patch. Tab. 5 compares the robustness performance of various methods under the YOLOv2 glass-box setting. As shown, our method achieves an ASR of $84.20\%$, far surpassing the second-best method, T-SEA, which records $36.15\%$. Moreover, AdvReal attains the highest Precision, Recall, and F1-Score ($14.87\%$, $15.80\%$, and $15.32\%$, respectively), demonstrating its superior ability to degrade detection performance even under occlusion. In contrast, other methods—such as AdvPatch, AdvTshirt, and AdvTexture—register ASRs below $30.00\%$ and exhibit considerably lower Precision, Recall, and F1-Scores. Notably, NatPatch achieves an ASR of only $3.46\%$, likely because it prioritizes natural appearance over adversarial potency. Overall, these results confirm the superior robustness and practical applicability of AdvReal as an adversarial patch.
Table 6: Transferability of different attack methods in YOLOv8, YOLOv11, and YOLOv12
@@
Figure 6: ASRs with different method and IOU thresholds in digital world. The glass-box is Faster RCNN and the confidence threshold is set as $0.5$.
@@ 4.5.2 Transfer performance across confidence threshold
Figure 7: ASRs with different detectors and confidence thresholds in digital world.
@@ 4.5.3 Mobility on advanced detectors
To further verify the transferability of AdvReal to the latest detectors, we evaluated various adversarial attack methods on three state-of-the-art (SOTA) detectors: YOLOv8 (2023s), YOLOv11 (2024s), and YOLOv12 (2025s). As shown in Tab. 6, the results indicate that AdvReal consistently produces significantly lower Precision, Recall, and F1-Scores across all detectors, underscoring its superior effectiveness in degrading object detection performance. Specifically, on YOLOv8, AdvReal achieved Precision, Recall, and F1-Score values of $56.55\%$, $32.68\%$, and $70.26\%$, respectively, outperforming all other methods. Similarly, on YOLOv11 and YOLOv12, AdvReal demonstrated outstanding performance, attaining Precision values of $66.37\%$ and $82.63\%$, Recall values of $32.47\%$ and $29.87\%$, and F1-Scores of $43.60\%$ and $43.88\%$, respectively. In contrast, methods such as AdvPatch, AdvTshirt, and AdvTexture performed substantially worse, particularly in Recall and F1-Score metrics. Notably, although AdvCaT performed reasonably well in Precision and F1-Score, its Recall value was excessively high (nearly $99\%$), indicating a weaker disruptive capability against the detection models. Overall, AdvReal demonstrated superior transferability across different detectors, confirming its robustness and generalization for cross-model adversarial attacks.
Figure 8: ASRs of adversarial patches trained with different glass-box detectors at various angles.
@@ 4.6.3 Visualization and Interpretability
We employed Grad-CAM Selvaraju et al. (2017) to generate activation maps that highlight the regions and salient features the detector attends to during its decision-making process. This helps reveal how AdvReal effectively deceives the object detection system by interfering with its attention. Fig. 9(a) and (b) present visualizations under good and poor lighting conditions in the digital space. In the non-adversarial pedestrian image on the left, the detector’s activation over the target region is smooth and inconspicuous. However, after introducing the adversarial patch, the detector’s attention clearly shifts toward the patch area. A similar trend is observed under poor lighting conditions: the pedestrian receives slightly partial attention, the adversarial patch in the corresponding image still dominates the activation map. In the physical world (Fig. 9(c) and (d)), the detection results and attention distributions at different distances further illustrate that the detector’s focus shifts from the entire pedestrian to the adversarial patch.
Fig. 10(a) shows a visualization of the detection box under conditions where $\text{IoU}=0.1$ and the confidence score is $0.1$. In this case, the detector identifies the patch itself as a person with low confidence, which aligns with the observed shift in attention toward the patch. This observation indicates a fundamental change in the detector’s decision-making basis caused by the patch. As shown in Fig. 10(b), when confronting the state-of-the-art detector YOLOv12, AdvReal still successfully misleads the model or significantly suppresses its confidence more effectively than existing adversarial methods, further validating its robustness and transferability.
@@ 5.1.2 Why 2D–3D joint optimization is effective?
5.1.3 Discoveries and recommendations of AI-safety in object detection system for AVs
The results reveals that the vulnerability of object detectors to physical adversarial patches varies considerably across different architectures. Single-stage detectors such as YOLOv11 and YOLOv12 are more to adversarial manipulation, while the earlier YOLOv8 version demonstrates higher robustness (Tab. 2 and Fig. 6). Two-stage detectors like Faster R-CNN exhibit greater resistance to physical attacks, likely due to its huge model parameters and intermediate region proposal mechanism that helps filter out adversarial cues Ren et al. (2016b). Meanwhile, transformer-based models such as D-DETR show even stronger robustness under multi-view conditions, maintaining stable performance across varying viewpoints (Fig. 8).
In the context of autonomous driving, image perception systems must strike a balance between real-time processing and resilience to adversarial threats. The experiments indicate that adversarial patches can significantly redirect the detector’s attention away from semantically meaningful regions, resulting in missed detections (Fig. 9 and Fig. 10). To mitigate these risks, we recommend reinforcing the structural robustness of detection models. Multi-stage and transformer-based architectures exhibit stronger spatial awareness and adversarial resistance. Moreover, detection heads should be equipped with mechanisms such as attention stabilization or feature denoising modules to counteract patch-induced interference. In summary, the development of safety-aware detectors requires not only stronger model architectures but also defense strategies grounded in real-world adversarial behavior.
@@ 6 Conclusion and Future works
This study proposed AdvReal, a robust adversarial patch generation framework designed to address challenges in physical-world attacks on object detection systems, including intra-class diversity, multi-angle variations, and complex environmental conditions. By incorporating 2D–3D joint optimization, physics-driven non-rigid surface modeling, and a realistic matching mechanism for lighting and geometry, AdvReal effectively bridges the gap between digital simulations and physical scenarios, thereby enhancing the transferability and robustness of adversarial patches. Experimental results across diverse object detectors and real-world settings demonstrate that AdvReal achieves state-of-the-art attack success rates, superior resilience to occlusion and lighting variations, and notable efficacy in redirecting detector attention. The proposed framework underscores the critical role of realistic 3D modeling and joint training in overcoming the limitations of conventional methods, providing a foundational benchmark for evaluating the adversarial robustness of AVs perception systems.
Looking ahead, several promising research directions warrant exploration to further enhance the framework and address current limitations. First, improving the generalization capability of adversarial patches across heterogeneous detector architectures requires consideration of structural differences in feature extraction and decision-making processes. Second, balancing adversarial effectiveness with visual naturalness could lead to more covert and practical attacks. Lastly, insights derived from AdvReal may inform the design of robust defense mechanisms, such as attention-stabilized detection modules or adversarial patch-aware training, to strengthen the resilience of intelligent object detection systems.

### v2

@@ 4.3 Ablation Study
Tab. 5 presents ablation experiments that systematically examine the contributions of three modules in the AdvReal framework—namely, non-rigid surfaces, realistic matching, and ShakeDrop. The results reveal that each module significantly influences both the attack success rate (ASR) and the perceptual quality, measured by average confidence (AC), of adversarial patches. Specifically, the exclusive use of the non-rigid surfaces module boosts ASR from $87.45\%$ to $91.56\%$ while simultaneously lowering AC from $59.68\%$ to $55.64\%$, thereby enhancing both attack efficacy and visual quality. Similarly, when applied individually, both the realistic matching and ShakeDrop modules improve ASR and reduce AC. Notably, the realistic matching module yields a particularly substantial improvement, raising ASR to $90.48\%$ and reducing AC to $55.39\%$.
Interestingly, the combined application of non-rigid surfaces and ShakeDrop results in a slight ASR decrease to $87.01\%$, yet it substantially improves perceptual quality, as evidenced by an AC reduction to $54.28\%$, indicating a synergistic benefit. Ultimately, the full integration of all three modules delivers the best performance, achieving an ASR of $93.72\%$ and an AC of $53.17\%$. The results confirm the cumulative impact of each component and highlight the framework’s ability to balance attack strength with visual imperceptibility.
@@ 4.4 Robustness
When the adversarial patch is partially occluded, its attack efficacy is markedly reduced. To evaluate the robustness of different algorithms under occlusion, we superimpose a $100\times 100$ pixel gray square at the center of a $300\times 300$ adversarial patch. Tab. 6 compares the robustness performance of various methods under the YOLOv2 glass-box setting. As shown in the table, our method achieves an ASR of $84.20\%$, significantly outperforming the second-best method, T-SEA, which records $36.15\%$. Moreover, AdvReal attains the highest values across all detection metrics—Precision, Recall, and F1-Score ($14.87\%$, $15.80\%$, and $15.32\%$, respectively)—demonstrating its strong ability to disrupt detection even under partial occlusion. In contrast, other methods such as AdvPatch, AdvTshirt, and AdvTexture record ASRs below $30.00\%$ and yield substantially lower detection metrics. NatPatch, in particular, achieves only a $3.46\%$ ASR, likely due to its prioritization of visual naturalness over adversarial strength.
Table 7: Transferability of different attack methods in YOLOv8, YOLOv11, and YOLOv12, the glass-box is YOLOv2
@@
Figure 6: ASRs with different method and IOU thresholds in digital world. The confidence threshold is set as $0.5$.
@@ 4.5.2 Transfer performance across confidence threshold
Figure 7: ASRs with different detectors and confidence thresholds in digital world. The glass-box detector used for adversarial training is marked on the top of each sub-image, and lines of different colors represent the corresponding detectors used for evaluation.
@@ 4.5.3 Mobility on advanced detectors
To further verify the transferability of AdvReal to the latest detectors, we evaluated various adversarial attack methods on three state-of-the-art (SOTA) detectors: YOLOv8 (2023s), YOLOv11 (2024s), and YOLOv12 (2025s). As shown in Tab. 7, the results indicate that AdvReal consistently yields significantly lower Precision, Recall, and F1-Scores across all detectors, highlighting its superior effectiveness in degrading object detection performance. Specifically, on YOLOv8, AdvReal achieved Precision, Recall, and F1-Score values of 56.55%, 32.68%, and 70.26%, respectively, outperforming all other methods. Similarly, on YOLOv11 and YOLOv12, AdvReal demonstrated outstanding performance, attaining Precision values of 66.37% and 82.63%, Recall values of 32.47% and 29.87%, and F1-Scores of 43.60% and 43.88%, respectively. In contrast, methods such as AdvPatch, AdvTshirt, and AdvTexture performed substantially worse, particularly in the Recall and F1-Score metrics. Notably, although AdvCaT performed reasonably well in Precision and F1-Score, its Recall value was exceptionally high (nearly 99%), indicating a weaker disruptive capability against the detection models. Overall, AdvReal exhibited superior transferability across different detectors, confirming its robustness and generalization in cross-model adversarial attacks.
4.6 Visual Quality and Detectability
4.6.1 Visual quality
To quantitatively evaluate the visual quality of different adversarial patches, we employ the PaQ-2-PiQ framework [58], a no-reference image quality prediction model trained on a large-scale database of real-world distortions and human judgments. This model outputs visual quality scores that are highly correlated with subjective perception, providing a reliable metric for assessing patch naturalness.
Table 8: Comparison of patch visual quality using PaQ-2-PiQ
AdvPatch AdvTexture AdvTshirt NatPatch TSEA AdvReal (Ours)
Visual Score $\uparrow$ 67.37 76.93 72.36 62.15 57.91 61.94
Category Good Good Good Fair Poor Fair
As summarized in Tab. 8, we evaluate six representative patch types and categorize them based on their perceptual scores. AdvTexture, AdvPatch, and AdvTshirt receive the highest scores and are rated as “Good”, reflecting their visually pleasing appearance or strong integration into background textures. In contrast, T-SEA and NatPatch obtain relatively low scores due to unnatural surface appearance or camouflage overuse.
Our method, AdvReal, achieves a visual quality score of 61.94 and is classified as “Fair”. While this reflects only a moderate level of visual realism, it is a direct result of prioritizing adversarial effectiveness under physical constraints. The optimization of texture alignment, relighting, and geometric adaptation inherently introduces visual artifacts that may reduce perceptual smoothness. Nonetheless, the generated patches remain physically plausible and deployable, striking a practical balance between attack strength and visual subtlety.
4.6.2 Detectability
To assess the detectability of adversarial patches from a visual security perspective, we employ NAPGuard [59], a dedicated framework for identifying naturalistic adversarial patches. NAPGuard distinguishes between aggressive and natural feature components within the patch and applies feature alignment and suppression to robustly detect adversarial content under visual camouflage. We evaluate all patch types using its pretrained model and report detection performance based on the AP@0.5 metric, which measures the ability to correctly localize patch regions.
Table 9: Patch detectability results using NAPGuard (AP@0.5)
AdvPatch AdvTexture AdvTshirt NatPatch TSEA AdvReal (Ours)
Average Confidence $\downarrow$ 0.6407 0.5952 0.4654 0.1147 0.8225 0.6212
Detection Rate $\downarrow$ 65.02% 63.07% 60.51% 61.37% 71.20% 60.67%
Tab. 9 summarizes the detection results across six representative methods. Patches with high adversarial saliency and minimal visual blending, such as T-SEA, exhibit the highest detectability ($0.8225$). In contrast, NatPatch, which emphasizes visual camouflage, has the lowest detection rate ($0.1147$). AdvReal achieves a moderate detectability score of $0.6212$, which is comparable to AdvPatch ($0.6407$) and AdvTexture ($0.5952$). This reflects the fact that while AdvReal is not specifically designed to evade detection, it also avoids conspicuous or structurally repetitive textures that would make it easily identifiable.
Fig. 8 provides the qualitative examples of patch detection results by NAPGuard. It can be seen that NAPGuard effectively localizes regions in patches with strong adversarial structure, while visually incoherent or low-signal patches evade detection. NatPatch typically triggers minimal activation due to its camouflage-like appearance, whereas T-SEA produces dense and easily localized responses, reflecting its high adversarial saliency but low visual subtlety. AdvReal exhibits consistent but non-dominant activation, suggesting that its adversarial patterns are strong enough to be partially detected, yet not overly distinguishable.
Figure 8: NAPGuard detection visualizations. Patches with aggressive features yield stronger activation. AdvReal maintains moderate detectability.
@@ 4.7 Physical Adversarial Attack Experiment
Figure 9: ASRs of adversarial patches trained with different glass-box detectors at various angles. All experimental results are obtained by adversarial training and evaluation on the models annotated in the subgraphs.
@@ 4.7.2 Different distances attacks
Figure 11: Localized attention shift visualized by Grad-CAM with zoom-in regions.
@@ 4.7.3 Visualization and Interpretability
We employed Grad-CAM [28] to generate class-discriminative activation maps that reveal the regions the detector attends to when performing person detection. The visualizations provide insight into how adversarial patches disrupt the detector’s semantic attention mechanisms, thereby influencing its decision-making process.
Fig. 10(a) and (b) present visualizations under good and poor lighting conditions in the digital domain. In the non-adversarial image, the detector’s attention is broadly distributed across the pedestrian’s torso and head, aligning with meaningful semantic regions. However, after introducing the adversarial patch, the attention distinctly shifts to the patch region. Fig. 11 (a) demonstrates that the adversarial patch significantly redirects the detector’s focus from the full-body region to patch-dominated areas. Fig. 11 (b) shows that despite changes in lighting, the patch consistently captures strong attention responses. Red boxes indicate zoomed-in views of the attended regions. Even under poor illumination, the patch continues to dominate the heatmap, suggesting robustness against lighting variation.
In the physical-world results shown in Fig. 10(c) and (d), similar behavior is observed at different distances. The detector’s attention weakens around the pedestrian body and intensifies around the patch location, especially when the patch is closer to the camera. This attention hijacking effect illustrates how AdvReal maintains adversarial influence under realistic constraints.
Fig. 12(a) and (b) compare detection responses across detectors under varying confidence and IoU thresholds. As the detector architecture becomes more advanced, its reliance on semantically aligned features increases, making attention shifts caused by the patch more semantically disruptive. Despite this, AdvReal maintains its effectiveness by consistently redirecting attention in all cases.
We also observe that the generated patches often contain structured gradients, edge-like formations, or repetitive textures. These features may partially resemble elements of the human form (e.g., limbs, contours), allowing the patch to falsely activate neuron clusters that normally respond to human-like regions. This suggests that AdvReal does not merely add noise, but introduces targeted feature-level deception.
The observations not only confirm the effectiveness of the attack but also offer interpretability of how the patch operates at the attention level. Such understanding could potentially inform future defense strategies—e.g., attention consistency regularization or saliency-guided anomaly suppression—to mitigate patch-based adversarial threats.
@@ 5.1.2 Why 2D–3D joint optimization is effective?
5.1.3 Discoveries and recommendations of AI-security in object detection system for AVs
The results reveal that the vulnerability of object detectors to physical adversarial patches varies considerably across different architectures. Single-stage detectors such as YOLOv11 and YOLOv12 are more susceptible to adversarial manipulation, while the earlier YOLOv8 version demonstrates higher robustness (Tab. 3 and Fig. 6). Two-stage detectors like Faster R-CNN exhibit greater resistance to physical attacks, likely due to their larger number of parameters and the presence of an intermediate region proposal mechanism that helps filter out adversarial cues [26]. Meanwhile, transformer-based models such as D-DETR show even stronger robustness under multi-view conditions, maintaining stable performance across varying viewpoints (Fig. 9).
The differences in adversarial vulnerability can be explained by the architectural nature of the detectors. One-stage models like YOLOv2 and YOLOv3 lack intermediate proposal filtering mechanisms and generate dense predictions directly from feature maps, making them especially sensitive to localized perturbations such as adversarial patches [55]. In contrast, two-stage detectors (e.g., Faster R-CNN) use a region proposal network to first identify high-confidence candidate regions, which acts as a form of internal filtering that helps suppress adversarial noise [56]. Transformer-based detectors (e.g., D-DETR) employ global self-attention mechanisms that further dilute the effect of spatially localized patches [57]. These structural differences account for the ASR variations reported in Tab. 3 and offer valuable insight for designing security-aware detectors in AVs.
In the context of autonomous driving, image perception systems must strike a balance between real-time processing and resilience to adversarial threats. The experiments indicate that adversarial patches can significantly redirect the detector’s attention away from semantically meaningful regions, resulting in missed detections (Fig. 10 and Fig. 12). To mitigate these risks, we recommend reinforcing the structural robustness of detection models. Architectures with multi-stage processing and transformer-based designs demonstrate stronger spatial contextual modeling and higher resistance to adversarial interference. Moreover, detection heads should be equipped with mechanisms such as attention stabilization modules or feature-level denoising blocks to counteract patch-induced interference. In summary, the development of security-aware detectors requires not only stronger model architectures but also defense strategies informed by empirical adversarial behavior in realistic environments.
5.1.4 Toward Identifying Malicious Attacks in AV Perception
Although adversarial attacks and natural detection errors may both lead to failed recognition, their underlying characteristics differ in systematic ways that can support intent-level distinction. Based on the and detectability in Sec. 4.6.2 and visualizations experiments in Sec. 4.7.3 , we observe that adversarial patches consistently redirect model attention to semantically irrelevant regions, whereas natural false positives usually result from ambiguous object boundaries or poor illumination without attention convergence. Furthermore, detection-based defense tools such as NAPGuard show higher response activations on structurally aggressive patterns, which are typically absent in naturally missed detections. These findings suggest that malicious attacks often exhibit spatial consistency, saliency concentration, and abnormal high resolution texture distribution. Thus, distinguishing adversarial attacks from natural false is feasible through joint analysis of attention behavior, patch localization, and detectability signals.
@@ 6 Conclusion and Future Works
This study proposed AdvReal, a robust adversarial patch generation framework designed to address challenges in physical-world attacks on object detection systems, including intra-class diversity, multi-angle variations, and complex environmental conditions. By incorporating 2D–3D joint optimization, physics-driven non-rigid surfaces modeling, and a realistic matching mechanism for lighting and geometry, AdvReal effectively bridges the gap between digital simulations and physical scenarios, thereby enhancing the transferability and robustness of adversarial patches. Experimental results across diverse object detectors and real-world settings demonstrate that AdvReal achieves state-of-the-art attack success rates, superior resilience to occlusion and lighting variations, and notable efficacy in redirecting detector attention. Beyond attack performance, this study also incorporates adversarial detectability evaluation as a potential defense perspective and discusses architectural strategies such as attention stabilization to improve model robustness. The proposed framework underscores the critical role of realistic 3D modeling and joint training in overcoming the limitations of conventional methods, providing a foundational benchmark for evaluating the adversarial robustness of AVs perception systems.
Looking ahead, several promising research directions warrant exploration to further enhance the framework and address current limitations. First, improving the generalization capability of adversarial patches across heterogeneous detector architectures requires consideration of structural differences in feature extraction and decision-making processes. Second, balancing adversarial effectiveness with visual naturalness could lead to more covert and practical attacks. Lastly, insights derived from AdvReal may inform the design of robust defense mechanisms. In particular, integrating adversarial training with patch-aware detection modules or developing co-evolution frameworks for joint attack-defense optimization offers promising potential for enhancing the security of real-world AV perception systems.
7 Acknowledgments
This work was supported by the fundamental research funds for the central universities.